# Simultaneous Task Allocation and Planning for Multi-Robots under Hierarchical Temporal Logic Specifications

### v3

@@ Abstract
Research in robotic planning with temporal logic specifications, such as syntactically co-safe Linear Temporal Logic (sc-LTL), has relied on single formulas. However, as task complexity increases, sc-LTL formulas become lengthy, making them difficult to interpret and generate, and straining the computational capacities of planners. To address this, we introduce a hierarchical structure to sc-LTL specifications with both syntax and semantics, proving it to be more expressive than flat counterparts. We conducted a user study that compared the flat sc-LTL with our hierarchical version and found that users could more easily comprehend complex tasks using the hierarchical structure. We develop a search-based approach to synthesize plans for multi-robot systems, achieving simultaneous task allocation and planning. This method approximates the search space by loosely interconnected sub-spaces, each corresponding to an sc-LTL specification. The search primarily focuses on a single sub-space, transitioning to another under conditions determined by the decomposition of automatons. We develop multiple heuristics to significantly expedite the search. Our theoretical analysis, conducted under mild assumptions, addresses completeness and optimality. Compared to existing methods used in various simulators for service tasks, our approach improves planning times while maintaining comparable solution quality.
@@ I Introduction
Formal methods, characterized by their mathematical rigor, are essential for specifying, developing, analyzing, and verifying software and hardware systems [3]. The recent trend of utilizing formal specifications, such as Linear Temporal Logic (LTL) [4], as high-level task specifications for multi-robots has garnered significant attention. This is because temporal logic specifications can encapsulate not just conventional point-to-point Boolean goals but also complex temporal requirements. While offering an expressive framework for task descriptions, temporal logic specifications also introduce computational challenges for solvers due to their complexity. Take LTL for an example. A typical method involves translating the formula into an automaton, a graphic representation. As shown in [5], which involves a task of collecting 5 keys and then opening 5 doors, it took approximately half an hour on a laptop with an Intel i7 CPU and 32 GB of RAM to convert the corresponding LTL into an automaton consisting of 65 nodes and 792 edges. Similarly, in our simulations, an automaton could not be generated within an hour on a laptop with Apple M2 Pro and 16G RAM. This issue stems from the usage of the “flat” form of LTL, that is, putting all requirements on robot behaviors into a single LTL formula, which tends to become difficult to solve for complex tasks. However, one key observation is that robot tasks often have a loose connection and can be broken down into smaller components. Studies suggest that humans prefer hierarchical task specification, which improves interpretability of planning and execution, making it easier to identify ongoing work and conveniently adjust unfeasible parts without affecting other components [6, 7]. Inspired by this, we propose a hierarchical form of LTL specifications.
Example 1
Consider a mobile manipulation task where robots are tasked with first picking and placing item $a$ and then fetching items $b$ and $c$ in any order. The “flat” specification is
[…]
The exact formulas can be found in Eq. (2), with syntax and semantics elaborated in Sec. IV. This representation effectively reduces the length of each formula, leading to smaller automatons and improving interpretability. $\square$
In this work, we first establish syntax rules for hierarchical syntactically co-safe LTL (sc-LTL) and then provide a framework for semantic interpretation. To do this, considering the context of multiple specifications, we introduce the concept of state-specification plan, which pairs the state of each robot with a specific specification it seeks to fulfill at a given time. Utilizing this concept, we devise an algorithm to check the satisfaction of a set of hierarchical sc-LTL specifications given a state-specification plan. Theoretically, we prove that the hierarchical form is more expressive compared to the flat form. Practically, we conducted a user study to show that using a hierarchical structure makes it easier to express complex tasks with temporal constraints. Next, we propose a planning algorithm with formal guarantees of completeness and optimality under mild assumptions. The bottom-up approach involves searching within a product graph that combines the environment and task space. This method’s advantage is it can simultaneously achieve task allocation and planning (STAP).
@@
Inspired by [8], we approximate the entire search space as a set of loosely interconnected sub-spaces, each corresponding to an individual sc-LTL specification in the hierarchical structure. The search process primarily occurs within a single sub-space, with transitions to adjacent sub-spaces happening under conditions determined by the decomposition of the automatons. Under certain assumptions regarding the state-specification plan, we demonstrate that our planning algorithm is sound, complete, and optimal. Additionally, we introduce several heuristics to expedite the search process, including minimizing the loose connections between sub-spaces and leveraging task-level progress to guide the search. We evaluate on extensive simulations on robot service tasks. Comparisons between our approaches with and without heuristics highlight a trade-off between computational time and solution quality. Our method shows significant improvements in computational time than work [8]. Notably, for complex tasks, [8] failed to produce solutions within a one-hour timeout. In scalability tests, our method demonstrated the ability to generate sub-optimal solutions in approximately three minutes for complex tasks involving up to 30 robots. Note that the flat version of these tasks could not be solved using graph-theoretic methods due to the inability to generate an automaton within a one-hour timeout. A glossary of key notations is provided in Tab. I.
@@
    We introduce a hierarchical form of LTL and prove it is more expressive than the flat form;
@@
    We conduct a user study to show that the hierarchical sc-LTL is easier to interpret compared to the flat sc-LTL.
@@
    We develop a search-based planning algorithm, achieving simultaneous task allocation and planning. This planner is, to the best of our knowledge, the first to offer both completeness and optimality for hierarchical sc-LTL;
@@ II-A Multi-robots Planning under Temporal Logic Specifications
As temporal logic formulas are used to tackle tasks involving multiple robots and complex environments, they inevitably become lengthy. Some attempts have been made to simplify this, such as merging multiple atomic propositions into one using logical operators [18, 21], or combining multiple sub-formulas into one formula using logical operators [14, 13, 26]. Despite these efforts, we still categorize these as the flat form. Notably, temporal operators, the feature distinguishing LTL from propositional logic, only appear on one side, either for the integration of merged propositions or inside the sub-formulas. Works [23, 26, 27, 28, 29, 30] have defined propositions involving more than one robot to encapsulate collaborative tasks. In our study, we introduce composite propositions that encompass more than one sub-formula and can be combined using temporal operators. Besides, various extensions of temporal logic have been introduced: [26] developed Counting Temporal Logic (CTL) to express the collective behavior of multiple robots. [31, 32] utilized Graph Temporal Logic (GTL) and Swarm Signal Temporal Logic (SwarmSTL) respectively to specify swarm characteristics like centroid positioning and density distribution. [27] introduced Capability Temporal Logic (CaTL) to encapsulate the diverse capabilities of robots. However, all works mentioned focus on using flat LTL specifications to describe the behaviors in multi-robot systems.
@@
The works most closely related to ours are [8, 39, 40, 41, 42]. Works [8, 39] approached STAP by breaking down temporal logic tasks into independent tasks, each of which can be completed by a robot. [40] built upon this idea, incorporating environmental uncertainties and failures into the planning. [41] expanded further into a multi-objective setting, integrating conflicting objectives such as cost and success probability into a Markov Decision Process (MDP). Our work presents key differences: (i) While the aforementioned studies focus on flat LTL specifications, we consider hierarchical specifications and can solve tasks which are unsolvable by these earlier methods. (ii) In the mentioned studies, tasks between robots are independent thus robot can execute their tasks in parallel. In contrast, our approach also assigns tasks with temporal dependencies to different robots, thus robots may depend on others. Note that [42], developed around the same time, is the only existing work that investigated decision-making under hierarchical sc-LTL. This work differs from ours in that, algorithmically, [42] proposed a hierarchical approach that first focused on task allocation and then planning, unlike STAP here. In terms of task decompoistion, [42] treated each transition in the automaton as an individual task to be allocated. In contrast, in this work individual tasks can consist of several transitions or even span the entire automaton, making each task potentially equivalent to multiple tasks as defined in [42]. Theoretically, the planning algorithm introduced in [42] lacks formal guarantees of either completeness or optimality.
@@ II-B Task Allocation and Planning
The domain of Multi-Robot Task Allocation (MRTA) is thoroughly studied, and various classifications of MRTA have been developed [1, 43], organizing the extensive research in this field. In terms of MRTA, the problem is this work can be is characterised by four aspects: (i) single-task robots (each robot is limited to executing one task at a time), (ii) single-robot tasks (only one robot is required for each task), (iii) instantaneous assignment (tasks are allocated under all available information at this moment), and (iv) cross-schedule interdependence (the assignment of a robot to a task is influenced not solely by its own schedule but also by the schedules of other robots within the system).
@@ II-C Hierarchical Task Models
Hierarchical reasoning enhances human understanding of the world [6, 7]. In classical AI planning, researchers crafted task models reflecting hierarchical structures by employing domain control knowledge [56]. These models have proven to be superior to flat models in terms of interpretability and efficiency, due to the significant reduction in the search space. Hierarchical Task Network (HTN) [57], a commonly used task model, exemplifies this. It presents a hierarchy of tasks, each of which can be executed if it’s primitive, or broken down into finer sub-tasks if it’s complex. Its plan comprises a set of primitive tasks applicable to the initial world state. Owing to its expressiveness, HTN has been implemented in the planning [58]. There are other hierarchical models such as AND/OR graphs [59] and sequential/parallel graphs [60].
Research combining hierarchical task models with LTL includes studies that use LTL to express temporally extended preferences over tasks and sub-tasks in HTN [61], as well as research into the expressive power of HTN in combination with LTL [62]. However, despite the widespread use of hierarchical task models in classical AI planning, it’s intriguing to note the lack of specification hierarchy in temporal logic robotic planning. Our work differs from these studies as we follow an inverse direction; instead of integrating LTL into HTN to express the goal of the planning problem, we incorporate HTN into LTL, allowing for hierarchical structures within multiple LTL formulas, making them more capable of expressing complex tasks than a single flat LTL formula.
@@
$\pi$ atomic propositions $\sigma$ infinite word over the alphabet
[…]
$\Diamond$ eventual operator ${\mathcal{Q}}_{\mathcal{B}}$ set of automaton states
$\square$ global operator ${\mathcal{Q}}_{\mathcal{B}}^{0}$ set of initial automaton states
$\mathcal{U}$ until operator $\Sigma=2^{{\mathcal{A}}{\mathcal{P}}}$ alphabet
$\phi$ specification $\to_{\mathcal{B}}$ transition relation in ${\mathcal{B}}$
${\mathcal{A}}{\mathcal{P}}$ set of atomic propositions ${\mathcal{Q}}_{\mathcal{B}}^{F}$ set of final automaton states
${\mathcal{T}}(r)$ transition system of robot $r$ $\rho_{\mathcal{B}}$ infinite run
${\mathcal{S}}_{r}$ states of robot $r$ ${\mathcal{L}}_{\mathcal{B}}$ accepted language of ${\mathcal{B}}$
@@
${\mathcal{G}}_{h}=({\mathcal{V}}_{h},{\mathcal{E}}_{h})$ specification hierarchy graph $c_{\mathcal{P}}$ cost function in ${\mathcal{P}}$
$\Phi^{k}$ set of specifications at the $k$-th level $\tau=\tau^{0}\ldots\tau^{h}$ state-specification plan
$\Phi^{K}$ set of specifications at the lowest level ${\mathcal{D}}_{\mathcal{B}}(\phi)$ decomposition set of specification $\phi$
$\Phi^{acc}$ set of satisfied specifications up to certain time ${\mathcal{P}}(\phi)$ product team model associated with specification $\phi$
@@ III Preliminaries
Notation: Let $\mathbb{N}$ denote the set of all integers, $[K]=\{1,\ldots,K\}$ denote the set of integers from 1 to $K$ and ||$\cdot$|| ⋅ | denote the cardinality of a set.
In this section, we formally describe Linear Temporal Logic (LTL) by presenting its syntax and semantics. Also, we briefly review preliminaries of automata-based LTL model checking.
@@
[…]
Definition III.1 (NBA)
[…]
  • ${\mathcal{Q}}_{{\mathcal{B}}}$ is the set of states;
  • ${\mathcal{Q}}_{{\mathcal{B}}}^{0}\subseteq{\mathcal{Q}}_{{\mathcal{B}}}$ is a set of initial states;
@@
  • $\rightarrow_{{\mathcal{B}}}\,\subseteq\,{\mathcal{Q}}_{{\mathcal{B}}}\times\Sigma\times{\mathcal{Q}}_{{\mathcal{B}}}$ is the transition relation;
  • ${\mathcal{Q}}_{{\mathcal{B}}}^{F}\subseteq{\mathcal{Q}}_{{\mathcal{B}}}$ is a set of accepting/final states.
An infinite run $\rho_{{\mathcal{B}}}$ of ${\mathcal{B}}$ over an infinite word $\sigma=\pi_{0}\pi_{1}\pi_{2}\dots$, $\pi_{k}\in\Sigma=2^{\mathcal{AP}}$, $\forall k\in\mathbb{N}$, is a sequence $\rho_{{\mathcal{B}}}=q_{{\mathcal{B}}}^{0}q_{{\mathcal{B}}}^{1}q_{{\mathcal{B}}}^{2}\dots$ such that $q_{{\mathcal{B}}}^{0}\in{\mathcal{Q}}_{{\mathcal{B}}}^{0}$ and $(q_{{\mathcal{B}}}^{k},\pi_{k},q_{{\mathcal{B}}}^{k+1})\in\rightarrow_{{\mathcal{B}}}$, $\forall k\in\mathbb{N}$. An infinite run $\rho_{{\mathcal{B}}}$ is called accepting if $\texttt{Inf}(\rho_{{\mathcal{B}}})\cap{\mathcal{Q}}_{{\mathcal{B}}}^{F}\neq\varnothing$, where $\texttt{Inf}(\rho_{{\mathcal{B}}})$ represents the set of states that appear in $\rho_{{\mathcal{B}}}$ infinitely often. The words $\sigma$ that produce an accepting run of ${\mathcal{B}}$ constitute the accepted language of ${\mathcal{B}}$, denoted by ${\mathcal{L}}_{{\mathcal{B}}}$. Then [63] proves that the accepted language of ${\mathcal{B}}$ is equivalent to the words of $\phi$, i.e., ${\mathcal{L}}_{{\mathcal{B}}}=\texttt{Words}(\phi)$.
The dynamics of robot $r$ is captured by a Transition System (TS) defined as follows:
@@
  • ${\mathcal{S}}_{r}=\{s_{r}\}$ is the set of discrete states of robot $r$;
  • $s_{r}^{0}$ is the initial state of robot $r$;

### v4

@@ Abstract
Research in robotic planning with temporal logic specifications, such as Linear Temporal Logic (LTL), has relied on single formulas. However, as task complexity increases, LTL formulas become lengthy, making them difficult to interpret and generate, and straining the computational capacities of planners. To address this, we introduce a hierarchical structure for a widely used specification type—LTL on finite traces (LTLf). The resulting language, termed H-LTLf, is defined with both its syntax and semantics. We further prove that H-LTLf is more expressive than its standard “flat” counterparts. Moreover, we conducted a user study that compared the standard LTLf with our hierarchical version and found that users could more easily comprehend complex tasks using the hierarchical structure. We develop a search-based approach to synthesize plans for multi-robot systems, achieving simultaneous task allocation and planning. This method approximates the search space by loosely interconnected sub-spaces, each corresponding to an LTLf specification. The search primarily focuses on a single sub-space, transitioning to another under conditions determined by the decomposition of automata. We develop multiple heuristics to significantly expedite the search. Our theoretical analysis, conducted under mild assumptions, addresses completeness and optimality. Compared to existing methods used in various simulators for service tasks, our approach improves planning times while maintaining comparable solution quality.
@@ I Introduction
Formal methods, characterized by their mathematical rigor, are essential for specifying, developing, analyzing, and verifying software and hardware systems [3]. The recent trend of utilizing formal specifications, such as Linear Temporal Logic (LTL) [4], as high-level task specifications for multi-robots has garnered significant attention. This is because temporal logic specifications can encapsulate not just conventional point-to-point Boolean goals but also complex temporal requirements. While offering an expressive framework for task descriptions, temporal logic specifications also introduce computational challenges for solvers due to their complexity. Take LTL for an example. A typical method involves translating the formula into an automaton, a graphic representation. As shown in [5], which involves a task of collecting 5 keys and then opening 5 doors, it took approximately half an hour on a laptop with an Intel i7 CPU and 32 GB of RAM to convert the corresponding LTL into an automaton consisting of 65 nodes and 792 edges. Similarly, in our simulations, an automaton could not be generated for a complex task within an hour on a laptop with Apple M2 Pro and 16G RAM. This issue stems from the usage of the standard “flat” form of LTL, that is, putting all requirements on robot behaviors into a single LTL formula, which tends to become difficult to solve for complex tasks. However, one key observation is that robot tasks often have a loose connection and can be broken down into smaller components. Studies suggest that humans prefer hierarchical task specification, which improves interpretability of planning and execution, making it easier to identify ongoing work and conveniently adjust unfeasible parts without affecting other components [6, 7]. Inspired by this, we propose a hierarchical form of a widely used specification type—LTL on finite traces (LTLf[8], which can be satisfied by finite sequence of system states. This characteristic makes LTLf apt for modeling and reasoning about systems with finite durations, such as those found in the robotics field.
@@
In this work, we first establish syntax and semantics for H-LTLf. To do this, considering the context of multiple specifications, we introduce the concept of state-specification sequence, which pairs the state of a system with a specification it seeks to fulfill at a given time. Theoretically, we prove that the hierarchical form is more expressive compared to the standard form. Practically, we conducted a user study to show that using a hierarchical structure makes it easier to express complex tasks with temporal constraints. Next, we propose a bottom-up planning algorithm that involves searching within a product graph that combines the environment and task space. This method’s advantage is it can simultaneously achieve task allocation and planning (STAP). Particularly, inspired by [9], we approximate the entire search space as a set of loosely interconnected sub-spaces, each corresponding to an individual LTLf specification in the hierarchical structure. The search process primarily occurs within a single sub-space, with transitions to adjacent sub-spaces happening under conditions determined by the decomposition of the automata. Under certain assumptions regarding the state-specification sequence, we demonstrate that our planning algorithm is sound, complete, and optimal. Additionally, we introduce several heuristics to expedite the search process, including minimizing the loose connections between sub-spaces and leveraging task-level progress to guide the search. We evaluate on extensive simulations on robot service tasks. Comparisons between our approaches with and without heuristics highlight a trade-off between computational time and solution quality. Our method shows significant improvements in computational time than work [9]. Notably, for complex tasks, [9] failed to produce solutions within a one-hour timeout. In scalability tests, our method demonstrated the ability to generate sub-optimal solutions in approximately three minutes for complex tasks involving up to 30 robots. The standard version of these tasks could not be solved using graph-theoretic methods due to the inability to generate an automaton within a one-hour timeout. A glossary of key notations is provided in Tab. I.
@@
    We introduce a hierarchical form of LTLf and prove it is more expressive than the standard form;
@@
    We conduct a user study to show that the H-LTLf  is easier to interpret compared to the standard LTLf;
@@
    We develop a search-based planning algorithm, achieving simultaneous task allocation and planning. This planner is, to the best of our knowledge, the first to offer both completeness and optimality for H-LTLf;
@@ II-A Multi-robots Planning under Temporal Logic Specifications
As temporal logic formulas are used to tackle tasks involving multiple robots and complex environments, they inevitably become lengthy. Some attempts have been made to simplify this, such as merging multiple atomic propositions into one using logical operators [19, 22], or combining multiple sub-formulas into one formula using logical operators [15, 14, 27]. Despite these efforts, we still categorize these as the standard form. Notably, temporal operators, the feature distinguishing LTL from propositional logic, only appear on one side, either for the integration of merged propositions or inside the sub-formulas. Works [24, 27, 28, 29, 30, 31] have defined propositions involving more than one robot to encapsulate collaborative tasks. In our study, we introduce composite propositions that encompass more than one sub-formula and can be combined using temporal operators. Besides, various extensions of temporal logic have been introduced: [27] developed Counting Linear Temporal Logic (cLTL) to express the collective behavior of multiple robots. [32, 33] utilized Graph Temporal Logic (GTL) and Swarm Signal Temporal Logic (SwarmSTL) respectively to specify swarm characteristics like centroid positioning and density distribution. [28] introduced Capability Temporal Logic (CaTL) to encapsulate the diverse capabilities of robots. However, all works mentioned focus on using standard LTL specifications to describe the behaviors in multi-robot systems.
@@
The works most closely related to ours are [9, 40, 41, 42, 43]. Works [9, 40] approached STAP by breaking down temporal logic tasks into independent tasks, each of which can be completed by a robot. [41] built upon this idea, incorporating environmental uncertainties and failures into the planning. [42] expanded further into a multi-objective setting, integrating conflicting objectives such as cost and success probability into a Markov Decision Process (MDP). Our work presents key differences: (i) While the aforementioned studies focus on standard LTL specifications, we consider hierarchical specifications and can solve tasks which are unsolvable by these earlier methods. (ii) In the mentioned studies, tasks between robots are independent thus robot can execute their tasks in parallel. In contrast, our approach also assigns tasks with temporal dependencies to different robots, thus robots may depend on others. Note that [43], developed around the same time, is the only existing work that investigated decision-making under H-LTLf. This work differs from ours in that, algorithmically, [43] proposed a hierarchical approach that first focused on task allocation and then planning, unlike STAP here. Theoretically, the planning algorithm introduced in [43] lacks formal guarantees of either completeness or optimality.
@@ II-B Task Allocation and Planning
The domain of Multi-Robot Task Allocation (MRTA) is thoroughly studied, and various classifications of MRTA have been developed [1, 44], organizing the extensive research in this field. In terms of MRTA, the problem is this work can be is characterized by four aspects: (i) single-task robots (each robot is limited to executing one task at a time), (ii) single-robot tasks (only one robot is required for each task), (iii) instantaneous assignment (tasks are allocated under all available information at this moment), and (iv) cross-schedule interdependence (the assignment of a robot to a task is influenced not solely by its own schedule but also by the schedules of other robots within the system).
@@ II-C Hierarchical Task Models
Studies [6, 7] suggest that hierarchical reasoning enhances humans’ ability to understand the world more effectively. In classical AI planning, researchers crafted task models reflecting hierarchical structures by employing domain control knowledge [57]. These models have proven to be superior to flat models in terms of interpretability and efficiency, due to the significant reduction in the search space. Hierarchical Task Network (HTN) [58], a commonly used task model, exemplifies this. It presents a hierarchy of tasks, each of which can be executed if it’s primitive, or broken down into finer sub-tasks if it’s complex. Its plan comprises a set of primitive tasks applicable to the initial world state. Owing to its expressiveness, HTN has been implemented in the planning [59]. There are other hierarchical models such as AND/OR graphs [60] and sequential/parallel graphs [61].
Research combining hierarchical task models with LTL includes studies that use LTL to express temporally extended preferences over tasks and sub-tasks in HTN [62], as well as research into the expressive power of HTN in combination with LTL [63]. However, despite the widespread use of hierarchical task models in classical AI planning, it’s intriguing to note the lack of specification hierarchy in temporal logic robotic planning. Our work differs from these studies as we follow an inverse direction; instead of integrating LTLf into HTN to express the goal of the planning problem, we incorporate HTN into LTLf, allowing for hierarchical structures within multiple LTLf formulas, making them more capable of expressing complex tasks than a single LTLf formula.
@@
$\sigma$ finite word over the alphabet ${\mathcal{A}}(\phi)$ NBA of specification $\phi$
${\mathcal{Q}}_{\mathcal{A}}^{0}$ set of initial automaton states ${\mathcal{Q}}_{\mathcal{A}}$ set of automaton states
$\Sigma=2^{{\mathcal{A}}{\mathcal{P}}}$ alphabet $\to_{\mathcal{A}}$ transition relation in ${\mathcal{A}}$
${\mathcal{A}}{\mathcal{P}}$ set of atomic propositions ${\mathcal{Q}}_{\mathcal{A}}^{F}$ set of final automaton states
${\mathcal{T}}(r)$ transition system of robot $r$ $\rho_{\mathcal{A}}$ finite run
${\mathcal{S}}_{r}$ states of robot $r$ ${\mathcal{L}}_{\mathcal{A}}$ accepted language of ${\mathcal{A}}$
@@
${\mathcal{G}}_{h}=({\mathcal{V}}_{h},{\mathcal{E}}_{h})$ specification hierarchy tree $c_{\mathcal{P}}$ cost function in ${\mathcal{P}}$
$\Phi_{k}$ set of specifications at the $k$-th level $\tau=\tau_{0}\ldots\tau_{h}$ state-specification sequence
$\Phi_{K}$ set of specifications at the lowest level ${\mathcal{D}}_{\mathcal{A}}(\phi)$ decomposition set of specification $\phi$
$\Phi_{\text{leaf}}$ set of leaf specifications ${\mathcal{P}}(\phi)$ product team model associated with specification $\phi$
@@ III Preliminaries
Notation: Let $\mathbb{N}$ denote the set of all integers, $[K]=\{0,\ldots,K\}$ and $[K]_{+}=\{1,\ldots,K\}$ represent the sets of integers from 0 to $K$ and from 1 to $K$, respectively, and ||$\cdot$|| ⋅ | denote the cardinality of a set.
In this section, we introduce LTLf by presenting its syntax and semantics, and automata-based LTLf model checking.
@@
LTLf [8] is a variant of LTL that is interpreted over finite sequences of states while maintaining the same syntax as standard LTL. A finite word $\sigma$ over the alphabet $2^{\mathcal{AP}}$ is defined as a finite sequence $\sigma=\sigma_{0}\sigma_{1}\ldots\sigma_{h}$ with $\sigma_{i}\in 2^{{\mathcal{A}}{\mathcal{P}}}$ for $i\in[h]$. The satisfaction of an LTLf formula $\phi$ over a sequence $\sigma$ at an instant $i$, for $i\in[h]$, is inductively defined as follows:
  • $\sigma,i\models\pi\;\;\text{iff}\;\pi\in\sigma_{i}$.
  • $\sigma,i\models\neg\phi\;\;\text{iff}\;\sigma,i\not\models\phi$.
  • $\sigma,i\models\phi_{1}\wedge\phi_{2}\;\;\text{iff}\;\sigma,i\models\phi_{1}\;\text{and}\;\sigma,i\models\phi_{2}$.
  • $\sigma,i\models\bigcirc\phi\;\;\text{iff}\;i<h\;\text{and}\;\sigma,i+1\models\phi$.
  • $\sigma,i\models\phi_{1}\,{\mathcal{U}}\,\phi_{2}$ iff for some $j$ such that $i\leq j\leq h$, we have that $\sigma,j\models\phi_{2}$, and for all $k,i\leq k<j$, we have that $\sigma,k\models\phi_{1}$.
A formula $\phi$ is satisfied by $\sigma$, denoted by $\sigma\models\phi$, if $\sigma,0\models\phi$. The language $\texttt{Words}(\phi)=\left\{\sigma|\sigma\models\phi\right\}$ is defined as the set of words that satisfy the formula $\phi$. An LTLf formula $\phi$ can be translated into a Nondeterministic Finite Automaton:
Definition III.1 (NFA)
A Nondeterministic Finite Automaton (NFA) ${\mathcal{A}}$ of an LTLf formula $\phi$ over $2^{\mathcal{AP}}$ is defined as a tuple ${\mathcal{A}}(\phi)=\left({\mathcal{Q}}_{{\mathcal{A}}},{\mathcal{Q}}_{{\mathcal{A}}}^{0},\Sigma,\rightarrow_{{\mathcal{A}}},\mathcal{Q}_{\mathcal{A}}^{F}\right)$, where
  • ${\mathcal{Q}}_{{\mathcal{A}}}$ is the set of states;
  • ${\mathcal{Q}}_{{\mathcal{A}}}^{0}\subseteq{\mathcal{Q}}_{{\mathcal{A}}}$ is a set of initial states;
@@
  • $\rightarrow_{{\mathcal{A}}}\,\subseteq\,{\mathcal{Q}}_{{\mathcal{A}}}\times\Sigma\times{\mathcal{Q}}_{{\mathcal{A}}}$ is the transition relation;
  • ${\mathcal{Q}}_{{\mathcal{A}}}^{F}\subseteq{\mathcal{Q}}_{{\mathcal{A}}}$ is a set of accepting/final states.
A finite run $\rho_{{\mathcal{A}}}$ of ${\mathcal{A}}$ over a finite word $\sigma=\sigma_{0}\sigma_{1}\dots\sigma_{h}$, $\sigma_{i}=2^{\mathcal{AP}}$, $\forall i\in[h]$, is a sequence $\rho_{{\mathcal{A}}}=q_{{\mathcal{A}}}^{0}q_{{\mathcal{A}}}^{1}\dots q_{{\mathcal{A}}}^{h+1}$ such that $q_{{\mathcal{A}}}^{0}\in{\mathcal{Q}}_{{\mathcal{A}}}^{0}$ and $(q_{{\mathcal{A}}}^{i},\sigma_{i},q_{{\mathcal{A}}}^{i+1})\in\rightarrow_{{\mathcal{A}}}$, $\forall i\in[h]$. A run $\rho_{{\mathcal{A}}}$ is called accepting if $q_{{\mathcal{A}}}^{h+1}\in{\mathcal{Q}}_{{\mathcal{A}}}^{F}$. The words $\sigma$ that produce an accepting run of ${\mathcal{A}}$ constitute the accepted language of ${\mathcal{A}}$, denoted by ${\mathcal{L}}_{{\mathcal{A}}}$. Then [64] proves that the accepted language of ${\mathcal{A}}$ is equivalent to the words of $\phi$, i.e., ${\mathcal{L}}_{{\mathcal{A}}}=\texttt{Words}(\phi)$.
The dynamics of robot $r$ is captured by a Transition System:
@@
  • ${\mathcal{S}}_{r}$ is the set of discrete states of robot $r$, and $s_{r}\in{\mathcal{S}}_{r}$ denotes a specific state;
  • $s_{r}^{0}\in{\mathcal{S}}_{r}$ is the initial state of robot $r$;
@@
A finite sequence of states $s_{r}^{0},s_{r}^{1},\ldots,s_{r}^{h}$ is referred to as a trace. Given the transition system ${\mathcal{T}}(r)$ of robot $r$ and the NFA ${\mathcal{A}}$ of an LTLf formula $\phi$, we can define the Product Automaton (PA) ${\mathcal{P}}(r,\phi)={\mathcal{T}}(r)\times{\mathcal{A}}(\phi)$ as follows [64]:
Definition III.3 (PA)
For a robot $r$ and an LTLf formula $\phi$, the Product Automaton is defined by the tuple ${\mathcal{P}}(r,\phi)=(\mathcal{Q}_{\mathcal{P}},\mathcal{Q}_{\mathcal{P}}^{0},\longrightarrow_{{\mathcal{P}}},\mathcal{Q}_{\mathcal{P}}^{F})$, where
  • $\mathcal{Q}_{\mathcal{P}}={\mathcal{S}}_{r}\times\mathcal{Q}_{{\mathcal{A}}}$ is the set of product states;
  • $\mathcal{Q}_{\mathcal{P}}^{0}=\{s_{r}^{0}\}\times\mathcal{Q}_{\mathcal{A}}^{0}$ is a set of initial states;
  • $\longrightarrow_{{\mathcal{P}}}\subseteq\mathcal{Q}_{\mathcal{P}}\times\mathcal{Q}_{\mathcal{P}}$ is the transition relation defined by the rule: $\frac{(s_{r}\rightarrow_{r}s^{\prime}_{r})\wedge(q\xrightarrow{{\mathcal{L}}_{r}\left(s_{r}\right)}_{{\mathcal{A}}}q^{\prime})}{q_{{\mathcal{P}}}=\left(s_{r},q\right)\longrightarrow_{\mathcal{P}}q_{{\mathcal{P}}}^{\prime}=\left(s^{\prime}_{r},q^{\prime}\right)}$. The transition from the state $q_{\mathcal{P}}\in\mathcal{Q}_{\mathcal{P}}$ to $q_{\mathcal{P}}^{\prime}\in\mathcal{Q}_{\mathcal{P}}$, is denoted by $(q_{\mathcal{P}},q_{\mathcal{P}}^{\prime})\in\longrightarrow_{\mathcal{P}}$, or $q_{\mathcal{P}}\longrightarrow_{\mathcal{P}}q_{\mathcal{P}}^{\prime}$;
  • $\mathcal{Q}_{\mathcal{P}}^{F}={\mathcal{S}}_{r}\times\mathcal{Q}_{\mathcal{A}}^{F}$ is a set of accepting/final states;
IV Hierarchical LTLf
We first present the syntax and semantics of H-LTLf, then we analyze its expressiveness.
IV-A Syntax of H-LTLf
Definition IV.1 (Hierarchical LTLf)
H-LTLf is structured into $K$ levels, labeled $L_{1},\ldots,L_{K}$, arranged from the highest to the lowest. Each level $L_{k}$ with $k\in[K]_{+}$ contains $n_{k}$ LTLf formulas. The H-LTLf specification is represented as $\Phi=\left\{\phi_{k}^{i}\,|\,k\in[K]_{+},i\in[n_{k}]_{+}\right\}$, where $\phi_{k}^{i}$ denotes the $i$-th LTLf formula at level $L_{k}$. Let $\Phi_{k}$ denote the set of formulas at level $L_{k}$, and let $\texttt{Prop}(\phi_{k}^{i})$ represent the set of propositions appearing in formula $\phi_{k}^{i}$. The H-LTLf follows these rules:
  1. 1.
    There is exactly one formula at the highest level: $n_{1}=1$.
  2. 2.
    Each formula at level $L_{k}$ consists either entirely of atomic propositions, i.e., $\texttt{Prop}(\phi_{k}^{i})\subseteq{\mathcal{A}}{\mathcal{P}}$, or entirely of formulas from the next lower level, i.e., $\texttt{Prop}(\phi_{k}^{i})\subseteq\Phi_{k+1}$.
  3. 3.
    Each formula at level $L_{k+1}$ appears in exactly one formula at the next higher level: $\phi_{k+1}^{i}\in\bigcup_{j\in[n_{k}]_{+}}\texttt{Prop}(\phi_{k}^{j})$ and $\texttt{Prop}(\phi_{k}^{j_{1}})\cap\texttt{Prop}(\phi_{k}^{j_{2}})=\varnothing$, for $j_{1},j_{2}\in[n_{k}]_{+}$ and $j_{1}\not=j_{2}$.
For a formula $\phi_{k}^{i}$ at a non-highest level, we slightly bend the notation to use $\phi_{k}^{i}$ to represent the same symbol at the higher level $L_{k-1}$, which we refer to as composite proposition. $\phi_{k}^{i}$ represents not only the $i$-th formula at level $L_{k}$, but also the corresponding composite proposition at level $L_{k-1}$. When $\phi_{k}^{i}$ appears at the right side in a certain formula, we consider it as a composite proposition. When it appears at the left side as a standalone formula, we refer to it as a specification.
Example 1 (H-LTLf)
Consider a task where robots are tasked with first picking and placing item $a$ and then fetching items $b$ and $c$ in any order. The LTLf specification is
where sub-formula $\Diamond(s_{x}\wedge\Diamond t_{x})$ denotes the event of first picking item $x$ from its source location $s_{x}$ and then placing it at the target location $t_{x}$. This formula is less interpretable by arranging all sub-formulas side by side. Furthermore, its automaton has 17 states and 102 transitions, which is unexpectedly large for a task of this complexity. The H-LTLf specifications of the same task are